OpenStudy (anonymous):

I made two graphs and found the correlation coefficients r^2. Help me with this question please?

OpenStudy (anonymous):

Is the difference in the two R2-values calculated in the two problems above significant? What do the two R2-values say about your cubic and quartic polynomial models? R^2 = 0.994408 for the first R^2 = 0.996725 for the second

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Morning! You know me by now. I'd like for you to explain in your own words what "R^2" means in practical terms, what its name is and how it's related to the correlation coefficient, r.

OpenStudy (anonymous):

Okay, I am all up for that.

OpenStudy (mathmale):

I'm feeling a little bad for having dumped such a big load of questions in your lap, but knowing what you know about r^2 will help me respond to your question.

OpenStudy (anonymous):

So far I just no that "r" is the correlation coefficient

OpenStudy (mathmale):

Right. We've talked about that a couple of times. Any clue at all what r^2 means?

OpenStudy (anonymous):

I would say that it is the correl. coeff. squared

OpenStudy (mathmale):

Certainly is. I'm curious: when in a situation in which you're not certain what something such as r^2 means, what do you typically do to investigate it? Have you tried any of those actions yet?

OpenStudy (anonymous):

No I have not tried anything of that sort. And sorry that it has been taking a while to reply. My mom and dad are both at work and both of my siblings have friends over, so I am the adult of the house today.

OpenStudy (mathmale):

Lot of responsibility for you, not to mention your having to learn these principles of statistics. r: correlation coefficient r^2: coefficient of determination. Have you heard or read anything about that?

OpenStudy (anonymous):

Not really, just correlation coefficient

OpenStudy (mathmale):

Even though I can explain the coefficient of determination, I'm always hoping that you'll look up such terms in whatever resources you have available: book, syllabus, Internet, etc. the coeff. of det., r^2, tells us what percentage of the variation in the dependent variable (usually y) is explained by the independent variable (usually x).

OpenStudy (mathmale):

You have in front of you two r^2 values. Both are very, very close to 1, aren't they? What this tells you is that your two different models very closely and accurately model the actual relationship between x and y in each case. If, however, r^2 were .6 (for example), only 60% of the variation in y would be explained by the model you've developed, meaning that the other 40% would be explained by factors other than x.

OpenStudy (mathmale):

Is the difference in the two R2-values calculated in the two problems above significant? What do the two R2-values say about your cubic and quartic polynomial models? R^2 = 0.994408 for the first R^2 = 0.996725 for the second Answering a question that has to do with how much one variable influences another quantity is at best an approximation, so I would respond to the first question, above, by saying something like, "No, the difference in the two r^2 values are not particularly significant. One rounds off to 0.99, the other to 1.00, a difference of only 1%. 2nd question: "These two r^2 values, rounding off as they do to 1.00 and 0.99 respectively, both indicate that the cubic and quartic models are excellent representations of the data on which they are founded: Just about 100% of the variation in y with x, in each case, is explained by the associated model." I don't expect you to come up with such fancy language, but would certainly hope that you could give a fairly accurate explanation, in your own words, of what r^2 stands for, how it's obtained and what it means.

OpenStudy (mathmale):

guess you're tied up with other responsibilities. Just respond when you're back. I'll be away from my computer for a while. MM